A rectangle and a triangle have equal areas. The length of the rectangle is 12 inches, and its width is 8 inches. If the base of the triangle is 32 inches, what is the length, in inches, of the altitude drawn to the base?
Answer by ewatrrr(24785)

 
Hi,
Area(rectangle) = l*w = 12in*8in = 96in^2
Area(triangle) = (1/2)32in*h = 96in^2 , h = 6in (h = 96in^2/16in)